40ml of ammonia gas taken in an eudiometer tube was subjected to sparks till the volume did not change any further .the volume was found to increase by 40ml.40ml of oxygen was then mixed and the mixture was exploded.the gases remained were 30ml.deduce the formula of ammonia

Grade:upto college level

2 Answers

Sunil Kumar FP
9 years ago
let he formula be NxHy
NxHy -----x/2 N2 + y/2H2
The equation is
x/2*40 +y/2*40=80
x+y=4 -----(1)

now N2 is left and h2 reacts with O2 to form water
H2 +1/2O2 ----H20
x/2*40 +40-10y=30
since volume left will be the combination of N2 and O2 left
2x-y=-1----(2)

solving 1 and 2 we get
x=1
y=3
thus the formula is NH3
Pramod Yadav
6 years ago
NxHy.....N2+H240ml. a ba+b=80......(1)H2 reacts with O2 only gives H2O...final product is N2, H2O and interacted O2... H2O being liquid get condensed....O2=(40-b/2) and N2=a...so from question;O2 + N2= 30.......(2).i.e a+40-b/2=30......(2)By solving 1 and 2 we get a=20 and b=60.Now by conserving N atom ...x ×40 = 2×20x= 1Again conserving H atomy × 40=2×60y=3Hence compound is NH3...